OpenStudy (anonymous):

z • 5 • y = y • 5 • z is an example of which algebraic property?

OpenStudy (anonymous):

Commutative property of multiplication

Parth (parthkohli):

This seems like Commutative.

Parth (parthkohli):

a * b * c = c * b * a

Parth (parthkohli):

Hell yea it is :P

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

i thought it was associative because it had the same numbers? lol of course im wrong lol

Parth (parthkohli):

No no no

Parth (parthkohli):

associative is like: a * (b * c) = (a * b) * c

Parth (parthkohli):

Heehee

OpenStudy (anonymous):

oohhh lol makes sense

OpenStudy (anonymous):

No assioctive has something to do with parenteses

OpenStudy (anonymous):

ohh

Parth (parthkohli):

Yep that's what I wrote.

Parth (parthkohli):

You can parenthesise any two numbers

OpenStudy (anonymous):

Lol it is really hard to reply on openstudy on an iPhone -_-

OpenStudy (anonymous):

lol its hard when your computer freezes everytime you start typing too lol